numbers are formed using all the digits 1, 2, 3, 4 ,5, 7, 8,9 without repetition. Determine the number of possible permutations in the following case. Show all steps in factorial form before final answers.

9 is before 1 but not necessarily adjacent to it. i.e, 912345678 or 234956178
Answer #1

1,2,3,4,5,6,7,8,9

we use permution for all arrangement

nPr=9P9=362880

=n!/ (n-r)!

=9!/ (9-9)!

=9!/1

=9*8*7*6*5*4*3*2*1=362880

but 9 is before 1 but not necessarily adjacent to it

8P8

nPr=8P8=40320

=n!/ (n-r)!

=8!/ (8-8)!

=8!/1

=8*7*6*5*4*3*2*1

=40320

but 9 is before 1 but not necessarily adjacent to it  40320

THANKS